OpenStudy (anonymous):

Linear approximation: f(x) = 1/(x^2+1) @ x = 1 approximate f(1.1)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

How did you derive f(x)?

OpenStudy (anonymous):

\[ f(x)=\frac1{x^2+1}\\ f'(x)=\frac{d}{dx} (x^2+1)^{-1}=-2x(x^2+1)^{-2}=-\frac{2x}{x^4+2x^2+1}\\ f(1.1)\simeq f(1)+f'(1)[1.1-1]\\ \ \ \ \ \ \ \ \ \ =\frac12-\frac12(0.1)\\ \ \ \ \ \ \ \ \ \ =\frac{10-1}{20}\\ \ \ \ \ \ \ \ \ \ =\frac{9}{20} \] Sorry, fixed.